give that y is directly proportionl to square of x and tha y = 16 when x= 2, calculate the value of y when x = 5
1
Expert's answer
2013-05-16T08:20:35-0400
Y is directly proportional to the square of x means thatthere is some number a, such that
y = a * x^2

Let's use the fact that y(2)=16 to find a:
16 = a * 2^2
16 = a * 4
a = 4

So, y = 4 * x^2

Therefore y(5) = 4 * 5^2 = 4 * 25 = 100

Answer: y(5)=100